# A Review of Hydrogen-Enabled Resilience Enhancement for Multi-Energy Systems

### v2

@@ II-C Security Issues of Hydrogen-related Technologies
Although resilience potential can be offered by hydrogen-related technologies in production, storage, transportation, and utilization, their security vulnerabilities also exist in each stage. During the production stage, especially in electrolysis and steam methane reforming, the risk of hydrogen leakage and explosion is significant due to high operating pressures and temperatures. During the storage stage, compressed hydrogen tanks face threats such as high-pressure rupture, while cryogenic liquid hydrogen presents risks of rapid venting. Long-term storage methods like salt caverns must address potential leakage and environmental contamination. During the transportation stage, hydrogen’s small molecular size increases the likelihood of leaks in pipelines or tube trailers, and mobile transport systems are vulnerable to traffic accidents or mechanical damage. Additionally, improper handling during transfer can lead to static ignition or pressure imbalances. During the utilization stage, poor ventilation in confined spaces (e.g., tunnels, vehicles, or buildings) can lead to hydrogen accumulation and fire hazards. Fuel cell systems may also face risks of overheating or malfunction if not properly managed. Across all stages, the growing reliance on automated control and digital monitoring introduces cybersecurity vulnerabilities[59], [60], e.g., hydrogen tank levels and hydrogen demand are vulnerable to false data injection attacks[60]. These security issues highlight the need for rigorous safety standards, advanced monitoring systems, and hydrogen-compatible materials to ensure safe deployment of hydrogen technologies.
@@
Index Challenges Hydrogen-related Example Formulas / Model Implications Formulations and Optimization Methods
@@
2 Mixed integer variables Investment locations of electrolyzers and hydrogen tanks $y_{i}^{\text{electrolyzer}}\in\{0,1\},\quad z_{j}^{\text{tank}}\in\{0,1\}$ MILP, MINLP, branch-and-bound, Lagrangian relaxation
@@
6 Nonlinear & nonconvex constraints Electrolyzer efficiency curve, pressure-flow relationships in pipelines $\eta^{\text{elec}}=f(P_{\text{input}}),\quad Q=k\cdot\sqrt{\Delta P}$ NLP, convex relaxation, piecewise linear approximation
7 Multi-energy coupling constraints Hydrogen enables electricity-gas-heat-transport sector integration $H_{t}^{\text{prod}}=\eta^{\text{elec}}\cdot P_{t}^{\text{elec}}/\omega^{\text{LHV}}$ Integrated MILP/MINLP, Benders decomposition
8 Multi-objective function Tradeoff between cost minimization and resilience/load recovery maximization $\min~\text{Cost}$; $\max~\text{Load~Recovery}$ MILP/MINLP, Multi-objective optimization/reinforcement learning
@@ II-D Optimization Challenges Related to Hydrogen-enabled Planning and Operation Enhancement
Integrating hydrogen into the planning and operation enhancement of MESs introduces a range of complex optimization challenges. Firstly, uncertainty in hydrogen demand complicates the optimization process, often requiring stochastic or robust optimization frameworks. Secondly, many hydrogen-related decisions, including infrastructure investments and on/off operational states of electrolyzers or compressors, involve binary or integer variables, resulting in mixed-integer programming problems. Thirdly, there are multi-timescale decision variables. Due to the physical limits on gas flow, thermal dynamics, and control systems, hourly updates of hydrogen storage level are sufficient for modeling storage dynamics. In contrast, batteries can be dispatched every 5 minutes to capture the fluctuations of renewable generation and load demand. When conducting operation optimization in hours, multi-timescale decision variables are introduced. Fourthly, hydrogen enables cross-temporal operations (e.g., producing and storing energy in low-demand periods for later use), which creates strong inter-temporal coupling and increases the dimensionality of the optimization problem. Fifthly, hydrogen transport via pipelines or trailers introduces spatial coupling, requiring network-based flow models with additional constraints. Sixthly, several key processes (e.g., electrolyzer efficiency, pressure-flow relationships in pipelines, and fuel cell performance) are inherently nonlinear, leading to nonconvex constraints that complicate solution tractability. Seventhly, the integration of hydrogen also deepens the coupling between electricity, gas, heat, and mobility sectors, leading to multi-energy coupling constraints. Finally, there are multi-objective functions in HMESs, e.g., system cost minimization and operation risk minimization, operation cost minimization and load recovery maximization. In summary, the above-mentioned challenges brought by the inclusion of hydrogen call for advanced solution methods, such as mixed-integer nonlinear programming, decomposition techniques, scenario-based planning, robust optimization[61], metaheuristics[60], convex relaxation, model predictive control[62], and deep reinforcement learning[63].
@@
For the ease of understanding, we summarize the optimization challenges introduced by adopting hydrogen in MES for resilience enhancement in Table IV, where $x$ denotes the first-stage decision variables (e.g., hydrogen storage or electrolyzer sizing); $\xi$ represents the uncertainty parameters (e.g., hydrogen demand, renewable availability); $\Xi$ is the uncertainty set; $c^{\top}x$ is the deterministic first-stage investment cost; $Q(x,\xi)$ is the recourse cost in the second stage given scenario $\xi$. $y_{i}^{\text{electrolyzer}}$ is a binary variable indicating whether an electrolyzer is installed at location $i$, $z_{j}^{\text{tank}}$ is a binary variable representing whether a hydrogen storage tank is constructed at node $j$, $y,z=1$ implies the component is built/active, while $y,z=0$ implies it is not. $H_{t}^{\text{stored}}$ and $B_{\tau}^{\text{stored}}$ denote the energy level in hydrogen tank and battery at time $t$ and $\tau$, respectively. $H_{t}^{\text{prod}}$ and $H_{t}^{\text{used}}$ denote the produced hydrogen and the consumed hydrogen at time $t$, respectively. $\mathcal{N}$ is the set of all nodes (e.g., cities, regions, or hydrogen hubs) in the hydrogen network, $\mathcal{T}$ is the set of time periods in the optimization horizon, $h_{ij,t}$ is the amount of hydrogen flow from node $i$ to node $j$ at time $t$, $h_{ji,t}$ is the amount of hydrogen flow from node $j$ to node $i$ at time $t$, $h^{\text{net}}_{i,t}$ is the net hydrogen injection at node $i$ at time $t$, which is positive if hydrogen is injected (e.g., from an electrolyzer) and negative if it is consumed (e.g., by a fuel cell or local demand). $\eta^{\text{elec}}$ is the efficiency of the electrolyzer, $P_{\text{input}}$ is the input electrical power, $f(\cdot)$ represents a nonlinear, possibly nonconvex function of power input, $Q$ is the hydrogen flow rate through a pipeline, $\Delta P$ is the pressure difference across the pipeline section, and $k$ is a physical constant related to pipe geometry and gas properties. $P_{t}^{\text{elec}}$ is the electrical power supplied to the electrolyzer at time $t$, $\omega^{\text{LHV}}$ denotes the lower heating value of hydrogen.
@@ III Hydrogen-enabled Resilience Enhancement Framework for MESs
As shown in Fig. 2, the framework of hydrogen-enabled resilience enhancement for MESs is provided, where four steps could be identified, i.e., resilience metrics, event-oriented contingency modeling, planning enhancement, and operation enhancement. To be specific, a resilience metric related to the investigated HMESs should be first defined, which would be incorporated into the objectives or constraints of planning or operation problem models. Then, according to the type of extreme events (e.g., hurricane, flood, ice disaster, earthquake, cyber attacks), the corresponding event models could be developed, which consist of event-related physical attributes[64], e.g., movement speed and landfall point of typhoons, and the intensity of earthquakes. Since such attributes depend on some unknown factors, the occurrence of an extreme event is full of uncertainties in the planning stage, e.g., there are many possible landfall scenarios when typhoons are considered[47]. After obtaining the physical attributes of extreme events, the stochastic impact of extreme events on the contingencies of MESs could be developed. Next, taking parameter uncertainty (e.g., renewable generation and multi-type loads), event uncertainty, and event-oriented contingency uncertainty into consideration, scenario generation could be conducted in preparation for planning and operation problems. In the planning stage, some permanent facilities will be deployed, e.g., SHSs, HESSs, hydrogen pipelines, and HRSs. During the operation stage, several kinds of responses can be made, e.g., preventive response, emergency response, and restoration response. During the preventive response, measures such as MHER location allocation, hydrogen procurement, and HESS pre-charging can be taken. During the emergency response, some measures are taken to minimize load shedding, including MHER scheduling, natural gas pipeline-assisted hydrogen use, HESS discharging, hydrogen demand adjustment, and hydrogen transformation. In the restoration stage, mobile hydrogen generation and storage, MHER routing and scheduling, HESS scheduling, hydrogen fuel cell scheduling, and hydrogen pipeline scheduling could be adopted to restore loads. By evaluating the practical system performance in the aspect of resilience, the rationality of the developed resilience metrics and event-oriented contingency modeling could be further improved.
@@ IV-A HMES Definitions and Architectures
Similar to the MES concept[65], an HMES is a system in which hydrogen, electricity, heat, cooling, gas, and transport interact optimally with each other at various levels to increase technical, economic, and environmental performance relative to “classical” energy systems whose sectors are treated separately. According to the spatial perspectives, HMESs can be divided into building-level, district-level, region-level, and nation-level. For simplicity, two typical architectures of HMESs are introduced, i.e., building-level HMESs and region-level HMESs. As shown in Fig. 3, the framework of a zero-carbon building-level HMES can be observed, and the corresponding practical system picture can be found in [66]. In Fig. 3, there are four kinds of energy flows, i.e., electricity flow, heat flow, cooling flow, and hydrogen flow. As far as electricity flow is concerned, the electricity generated from photovoltaic (PV) panels and the fuel cell is used to satisfy the electricity demand of an electric boiler, a ground-source heat pump (GSHP), an absorption chiller, an electrolyzer, and electric demand in a commercial building. As to heat flow, the heat demand is satisfied by the electric boiler, GSHP, fuel cell, and hot water tank. Similarly, cooling demand is satisfied by the GSHP, absorption chiller, and cold water tank in the cooling flow. As to hydrogen flow, PV generation can be used for the electrolyzer to produce hydrogen, which will be stored in the hydrogen tank for future use by driving fuel cells. When the hydrogen stored in the hydrogen tank is not enough, hydrogen purchasing from the market is required. In Fig. 4, the architecture of a region-level HMES can be observed[67], which consists of three types of communities, e.g., residential communities, commercial communities, and industrial communities. In residential and commercial communities, electricity/cold/heat/hydrogen demands can be identified. In industrial communities, electricity/hydrogen demands can be observed. All communities are connected through electricity transportation networks, hydrogen transmission pipelines, and transportation networks. With the help of transportation networks, MHERs (e.g., HTTs and HFCVs can be redistributed among multiple communities for improving network resilience, e.g., providing emergent power supply or hydrogen supply to a community when it suffers from an extreme event (e.g., power outages or hydrogen pipeline leakage).
@@
Figure 4: Typical HMES architecture (region-level)

### v3

@@ II-C Vulnerabilities of Hydrogen-related Technologies
Although resilience potential can be offered by hydrogen-related technologies in the chain of production, storage, transportation, and utilization, security vulnerabilities also exist in each stage, which can create new disturbance propagation paths and may weaken rather than strengthen system resilience under adverse conditions. To be specific, in the production stage, especially in electrolysis and steam methane reforming, the risk of hydrogen leakage and explosion is significant due to high operating pressures and temperatures. In the storage stage, compressed hydrogen tanks face threats such as high-pressure rupture, while cryogenic liquid hydrogen presents risks of rapid venting. Long-term storage methods like salt caverns must address potential leakage and environmental contamination. In the transportation stage, hydrogen’s small molecular size increases the likelihood of leaks in pipelines or tube trailers, and mobile transport systems are vulnerable to traffic accidents or mechanical damage. Additionally, improper handling during transfer can lead to static ignition or pressure imbalances. In the utilization stage, inadequate ventilation may cause hydrogen accumulation and fire hazards. Moreover, failures in electrolyzers, hydrogen storage tanks, hydrogen pipelines, or fuel cell systems may compromise local energy supply and even lead to load shedding. In addition, the growing digitalization of hydrogen infrastructures introduces cybersecurity threats[59], [60], such as false data injection attacks on tank levels and demand information[60]. Therefore, hydrogen should be regarded as a technology with both resilience benefits and resilience risks, and its positive contribution to MES resilience depends on proper design, operation, protection, and cybersecurity measures.
@@
Index Challenges Hydrogen-related Example Formulas / Model Implications Formulations and Optimization Methods/Techniques
@@
2 Mixed integer variables Investment locations of electrolyzers and hydrogen tanks $y_{i}^{\text{electrolyzer}}\in\{0,1\},\qquad z_{j}^{\text{tank}}\in\{0,1\}$ MILP, MINLP, branch-and-bound, Lagrangian relaxation
@@
6 Nonlinear & nonconvex constraints Load-dependent electrolyzer/fuel cell efficiencies, part-load operating regimes, nonlinear hydrogen pipeline hydraulics, compression/liquefaction energy consumption, and nonlinear hydrogen storage state relations $\eta_{t}^{\mathrm{EL}}=f(P_{t}^{\mathrm{EL}})$, $\eta_{t}^{\mathrm{FC}}$ = $g(P_{t}^{\mathrm{FC}})$, $H_{t}^{\mathrm{prod}}$ $\frac{\eta_{t}^{\mathrm{EL}}P_{t}^{\mathrm{EL}}\Delta t}{\omega^{\mathrm{L}}}$, $P_{t}^{\mathrm{FC,out}}$ $\frac{\eta_{t}^{\mathrm{FC}}H_{t}^{\mathrm{FC,in}}\omega^{\mathrm{L}}}{\Delta t}$, $h_{ij,t}$$k_{i,j}(p_{i,t},~p_{j,t})$, $E_{t}^{\mathrm{c}}$$\varphi(q_{t}^{\mathrm{H_{2}}},p_{t}^{\mathrm{in}},p_{t}^{\mathrm{out}})$$p_{t}^{\mathrm{tk}}$$\phi(m_{t}^{\mathrm{H_{2}}},T_{t},V_{t})$ MINLP, NLP, piecewise linear approximation, convex/cone relaxation, sequential linearization, McCormick envelopes, metaheuristics, and deep reinforcement learning
7 Multi-energy coupling constraints Hydrogen enables electricity-gas-heat-transport sector integration $P_{t}^{\mathrm{FC,e}}=\frac{\eta_{t}^{\text{FC,e}}H_{t}^{\mathrm{FC,in}}\omega^{\mathrm{L}}}{\Delta t}$,  $Q_{t}^{\mathrm{FC,h}}=\frac{\eta_{t}^{\text{FC,h}}H_{t}^{\mathrm{FC,in}}\omega^{\mathrm{L}}}{\Delta t}$ Integrated MILP/MINLP, Benders decomposition
8 Multi-objective functions Tradeoff between cost minimization and resilience/load recovery maximization $\min~\text{Cost}$; $\max~\text{Load~Recovery}$ MILP/MINLP, Multi-objective optimization/reinforcement learning
@@ II-D Optimization Challenges Related to Hydrogen-enabled Planning and Operation Enhancement
Integrating hydrogen into the resilience-oriented planning and operation of MESs introduces a range of complex optimization challenges. Firstly, uncertainty in hydrogen demand complicates the optimization process, often requiring stochastic or robust optimization frameworks. Secondly, many hydrogen-related decisions, including infrastructure investments and on/off operational states of electrolyzers or compressors, involve binary or integer variables, resulting in mixed-integer programming problems. Thirdly, there are multi-timescale decision variables. Due to the physical limits on gas flow, thermal dynamics, and control systems, hourly updates of hydrogen storage level are sufficient for modeling storage dynamics. In contrast, batteries could be dispatched every 5 minutes to capture the fluctuations of renewable generation and load demand. When conducting operation optimization on an hourly basis, multi-timescale decision variables would arise. Fourthly, hydrogen enables cross-temporal operations (e.g., producing and storing energy in low-demand periods for later use), which creates strong inter-temporal coupling and increases the dimensionality of the optimization problem. Fifthly, hydrogen transport via pipelines or trailers introduces spatial coupling, requiring network-based flow models with additional constraints. Sixthly, there are many nonlinear or even nonconvex constraints involved in the hydrogen production and the associated compression, storage, transport, and utilization processes, which complicate solution tractability. Taking electrolyzers, fuel cells, hydrogen pipelines, hydrogen compressors/liquefiers, and hydrogen storage tanks into consideration, there are nonlinear or nonconvex constraints incurred by load-dependent electrolyzer/fuel cell efficiencies, part-load operating regimes of electrolyzers/fuel cells, nonlinear hydrogen pipeline hydraulics, energy consumption of hydrogen compressors/liquefiers, and nonlinear hydrogen storage state relations. Seventhly, the integration of hydrogen also deepens the coupling between electricity, gas, heat, and mobility sectors, leading to multi-energy coupling constraints. Finally, there are multi-objective functions in HMESs, e.g., system cost minimization and operation risk minimization, operation cost minimization and load recovery maximization. It is worth noting that the mathematical models of hydrogen equipment do not necessarily introduce features that are entirely absent from all other MES components, since nonlinearities, mixed-integer decisions, and temporal couplings may also arise in other devices such as CCHP units, batteries, and gas networks. However, compared with conventional MES equipment, hydrogen-related devices tend to exhibit these characteristics in a more pronounced and integrated manner, especially when hydrogen production, compression, storage, transport, and utilization are modeled simultaneously. In order to address the above challenges, more advanced solution methods should be adopted, such as mixed-integer nonlinear programming, decomposition techniques, scenario-based planning, robust optimization[61], metaheuristics[60], convex relaxation, model predictive control[62], and deep reinforcement learning[63].
@@
For the ease of understanding, we summarize the optimization challenges introduced by adopting hydrogen in MESs for resilience enhancement in Table IV. In line 1, $x$ denotes the first-stage decision variables (e.g., hydrogen storage or electrolyzer sizing); $\xi$ represents the uncertainty parameters (e.g., hydrogen demand, renewable availability); $X$ is the uncertainty set; $c^{\top}x$ is the deterministic first-stage investment cost; $Q(x,\xi)$ is the recourse cost in the second stage given scenario $\xi$. In line 2, $y_{i}^{\text{electrolyzer}}$ is a binary variable indicating whether an electrolyzer is installed at location $i$, $z_{j}^{\text{tank}}$ is a binary variable representing whether a hydrogen storage tank is constructed at node $j$, $y,z=1$ implies the component is built/active, while $y,z=0$ implies it is not. In line 3, $H_{t}^{\text{stored}}$ and $B_{\tau}^{\text{stored}}$ denote the energy level in hydrogen tanks and batteries at time $t$ and $\tau$, respectively. In line 4, $H_{t}^{\text{prod}}$ and $H_{t}^{\text{used}}$ denote the produced hydrogen and the consumed hydrogen at time $t$, respectively. In line 5, $\mathcal{N}$ is the set of all nodes (e.g., cities, regions, or hydrogen hubs) in the hydrogen network, $\mathcal{T}$ is the set of time periods in the optimization horizon, $h_{ij,t}$ is the amount of hydrogen flow from node $i$ to node $j$ at time $t$, $h_{ji,t}$ is the amount of hydrogen flow from node $j$ to node $i$ at time $t$, $h^{\text{net}}_{i,t}$ is the net hydrogen injection at node $i$ at time $t$, which is positive if hydrogen is injected (e.g., from an electrolyzer) and negative if it is consumed (e.g., by a fuel cell or local demand). In line 6, $\eta_{t}^{\mathrm{EL}}$ and $\eta_{t}^{\mathrm{FC}}$ denote the electrolyzer and fuel cell efficiencies at time $t$, respectively; $f$, $g$, $k_{i,j}$, $\varphi$ and $\phi$ are nonlinear or even nonconvex functions. $P_{t}^{\mathrm{EL}}$ and $P_{t}^{\mathrm{FC}}$ denote the electrolyzer input power and fuel cell operating power, respectively; $\Delta t$ is the time-step duration; and $\omega^{\mathrm{L}}$ is the lower heating value of hydrogen. $P_{t}^{\mathrm{FC,out}}$ denotes the fuel cell electric output power, and $H_{t}^{\mathrm{FC,in}}$ is the hydrogen input to the fuel cell. $p_{i,t}$ and $p_{j,t}$ denote the nodal hydrogen pressures. $E_{t}^{\mathrm{c}}$ denotes the compressor energy consumption, $q_{t}^{\mathrm{H_{2}}}$ is the hydrogen flow through the compressor, and $p_{t}^{\mathrm{in}}$ and $p_{t}^{\mathrm{out}}$ denote the inlet and outlet pressures, respectively. $p_{t}^{\mathrm{tk}}$ is the hydrogen tank pressure, while $m_{t}^{\mathrm{H_{2}}}$, $T_{t}$, and $V_{t}$ denote the stored hydrogen mass, storage temperature, and tank volume at slot $t$, respectively. In line 7, $\eta_{t}^{\mathrm{FC,e}}$ and $\eta_{t}^{\mathrm{FC,h}}$ are the electrical and thermal conversion efficiencies of the fuel cell, respectively. $P_{t}^{\mathrm{FC,e}}$ and $Q_{t}^{\mathrm{FC,h}}$ denote the electrical and thermal outputs of the fuel cell, respectively.
@@ III Hydrogen-enabled Resilience Enhancement Framework for MESs
As shown in Fig. 2, the framework of hydrogen-enabled resilience enhancement for MESs is provided, where four steps could be identified, i.e., resilience metrics, event-oriented contingency modeling and uncertainty processing, planning enhancement, and operation enhancement. To be specific, a resilience metric related to the investigated HMESs should be first defined, which would be incorporated into the objectives or constraints of planning or operation problem models. Then, according to the type of extreme events (e.g., hurricane, flood, ice disaster, earthquake, and cyber attacks), the corresponding event models could be developed, which consist of event-related physical attributes[64], e.g., movement speed and landfall point of typhoons, and the intensity of earthquakes. Since such attributes depend on some unknown factors, the occurrence of an extreme event is full of uncertainties in the planning stage, e.g., there are many possible landfall scenarios when typhoons are considered[47]. After obtaining the physical attributes of extreme events, the contingencies of MESs incurred by extreme events need to be modeled. Typically, stochastic multi-type contingencies are considered in existing works[22]. In order to cope with the above-mentioned event uncertainties and contingency uncertainties, as well as parameter uncertainty (e.g., renewable generation and multi-type loads), uncertainty processing methods will be adopted, e.g., scenario-based methods, uncertainty-set-based methods, and ambiguity-set-based methods. Such methods can provide scenario parameter information, feasible decision space, and probability distribution sets for planning and operation optimization problems. More details about uncertainty processing methods and extreme-event scenario generation methods will be introduced in the Table VIII and Table X in Section VI, respectively. In the planning stage, some permanent facilities will be deployed, e.g., SHSs, HESSs, hydrogen pipelines, and HRSs. During the operation stage, several kinds of responses can be made, e.g., preventive response, emergency response, and restoration response. During the preventive response, measures such as MHER location allocation, hydrogen procurement, and HESS pre-charging can be taken. During the emergency response, some measures are taken to minimize load shedding, including MHER scheduling, natural gas pipeline-assisted hydrogen use, HESS discharging, hydrogen demand adjustment, and hydrogen transformation. In the restoration stage, mobile hydrogen generation and storage, MHER routing and scheduling, HESS scheduling, hydrogen fuel cell scheduling, and hydrogen pipeline scheduling could be adopted to restore loads. By evaluating the practical system performance in the aspect of resilience, the rationality of the developed resilience metrics and event-oriented contingency modeling could be further improved.
@@ IV-A HMES Definitions and Architectures
Similar to the MES concept[65], an HMES is a system in which hydrogen, electricity, heat, cooling, gas, and transport interact optimally with each other at various levels to increase technical, economic, and environmental performance relative to “classical” energy systems whose sectors are treated separately. According to the spatial perspectives, HMESs can be divided into building-level, district-level, region-level, and nation-level. For simplicity, two typical architectures of HMESs are introduced, i.e., building-level HMESs and region-level HMESs. As shown in Fig. 3, the framework of a zero-carbon building-level HMES can be observed, and the corresponding practical system picture can be found in [66]. In Fig. 3, there are four kinds of energy flows, i.e., electricity flow, heat flow, cooling flow, and hydrogen flow. As far as electricity flow is concerned, the electricity generated from photovoltaic (PV) panels and the fuel cell is used to satisfy the electricity demand of an electric boiler, a ground-source heat pump (GSHP), an absorption chiller, an electrolyzer, and electric demand in a commercial building. As to heat flow, the heat demand is satisfied by the electric boiler, GSHP, fuel cell, and hot water tank. Similarly, cooling demand is satisfied by the GSHP, absorption chiller, and cold water tank in the cooling flow. As to hydrogen flow, PV generation can be used for the electrolyzer to produce hydrogen, which will be stored in the hydrogen tank for future use by driving fuel cells. When the hydrogen stored in the hydrogen tank is not enough, hydrogen purchasing from the market is required. In Fig. 4, the architecture of a region-level HMES can be observed[67], which consists of three types of communities, e.g., residential communities, commercial communities, and industrial communities. In residential and commercial communities, electricity/cold/heat/hydrogen demands can be identified. In industrial communities, electricity/hydrogen demands can be observed. All communities are connected through electricity distribution networks, hydrogen transmission pipelines, and road transportation networks. With the help of transportation networks, MHERs (e.g., HTTs and HFCVs) can be redistributed among multiple communities for improving network resilience, e.g., providing emergent power supply or hydrogen supply to a community when it suffers from an extreme event (e.g., power outages or hydrogen pipeline leakage).
@@
Figure 4: Typical HMES architecture (region-level)[67]